### 30 Mar 2026 v2

# Topological bounds on the dynamical growth rate of chemical reaction networks

Abstract
Growth and decay are system-level properties of chemical reaction networks (CRNs) relevant from prebiotic chemistry to cellular metabolism. Their properties are typically analyzed through the kinetics of particular models, which requires specification of the full set of kinetic laws and parameters. In this work, we derive stoichiometry-based constraints on the growth (or shrinkage) rate, in the balanced-growth regime of scalable CRNs. The resulting bounds are controlled by a topological quantity, the maximum amplification factor, defined via a von Neumann max–min problem over feasible fluxes as illustrated by numerical tests on random-network ensembles of CRNs. We argue for the relevance of our results in the context of origin of life studies, but also for designing synthetic chemical reaction networks.
Autocatalysis and positive feedback are ubiquitous in biochemical organization and in candidate prebiotic chemistries: they enable self-maintenance, replication-like amplification, and, in open settings, persistent growth [eigen1977principle, baum2023ecology, ameta_self-reproduction_2021]. Stoichiometric frameworks have recently made the identification of autocatalytic subnetworks operational via purely structural criteria and algorithms that identify minimal autocatalytic motifs and their embeddings inside large networks [blokhuis2020universal, gagrani2024polyhedral, golnik2025enumeration, andersen_defining_2020]. This progress raises a natural quantitative question: once a network is topologically capable of amplification, what controls how fast it can grow or decay?
From a dynamical perspective, growth rates depend on nonlinear kinetics and on rate constants that can vary over orders of magnitude and are often poorly constrained. This creates a gap: topology-based notions of autocatalysis certify feasibility, while kinetics determines realized performance. Bridging this gap requires a mixed kinetic–topological metric that quantifies amplification strength. Prior structural results in mass-action settings exist in restricted regimes (e.g., special coupling or dilute limits) [despons_structural_2024, unterberger2022stoechiometric, nandan_p_autocatalytic_2025]; here we identify a broad balanced-growth regime in which a single metric controls the achievable growth rate.
Maximum Amplification Factor— A CRN with internal species set $\mathcal{S}$ and reaction set $\mathcal{R}$ is specified by nonnegative $|\mathcal{S}|\times|\mathcal{R}|$ output and input matrices $(\mathbb{S}^{+},\mathbb{S}^{-})$, where $|\mathcal{S}|$ and $|\mathcal{R}|$ are the number of species and reactions in the networks, respectively. The associated stoichiometric matrix is then given by $\mathbb{S}=\mathbb{S}^{+}-\mathbb{S}^{-}$. In this work, we do not assume reversible reactions and each column of the stoichiometric matrix corresponds to a unique one-way reaction. We consider the dynamics
$\dot{\mathbf{n}}=\mathbb{S}\,\mathbf{J}(\mathbf{n}),$ (1)
where $\mathbf{n}$ collects populations and $\mathbf{J}(\mathbf{n})$ denotes the one-way reaction fluxes. Independently of kinetics, we also consider nonnegative flux vectors $\mathbf{x}\geq 0$ as arbitrary reaction fluxes. Given $\mathbf{x}$, the induced species-wise input and output rates are $\mathbb{S}^{-}\mathbf{x}$ and $\mathbb{S}^{+}\mathbf{x}$, respectively (SM). The topological quantity introduced next is defined by optimizing over such fluxes.
We define the Amplification Factor (AF) of a flux vector $\mathbf{x}$ as the species-wise minimum ratio of output to input rates
$\alpha(\mathbf{x})\;=\;\;\min_{s\in\mathcal{S}}\frac{(\mathbb{S}^{+}\mathbf{x})_{s}}{(\mathbb{S}^{-}\mathbf{x})_{s}}.$ (2)
The Maximum Amplification Factor (MAF) is the maximal value of the amplification factor over feasible fluxes,
$\alpha(\mathbb{S}^{+},\mathbb{S}^{-})\,=\,\max_{\mathbf{x}>\bm{0}}\,\alpha(\mathbf{x})\,=\,\max_{\mathbf{x}>\bm{0}}\;\min_{s\in\mathcal{S}}\frac{(\mathbb{S}^{+}\mathbf{x})_{s}}{(\mathbb{S}^{-}\mathbf{x})_{s}},$ (3)
with the convention that species with $(\mathbb{S}^{-}\mathbf{x})_{s}=0$ do not constrain the minimum (they correspond to $\infty$). Note that $\mathbf{x}>\mathbf{0}$ denotes that $\mathbf{x}\geq\mathbf{0}$ but $\mathbf{x}\neq\mathbf{0}$. Inspired by von Neumann’s technological expansion rate in input-output models of economics ([neumann1945model], [sargent2019neumann], Sec. 9.5 in [gale1989theory]), the MAF was introduced for hypergraphs in [blanco2025identifyingselfamplifyinghypergraphstructures]. From now on, we will denote the MAF with $\alpha$ for clarity of presentation.
Autonomous and autocatalytic networks— We say that a CRN is autonomous if every reaction has at least one reactant and at least one product and every species appears as a reactant in at least one reaction and as a product in at least one reaction. In practice, reactions that effectively create or remove matter are modeled through chemostatted species representing exchange with an environment. Given an arbitrary network, an autonomous subnetwork can be obtained by iteratively removing chemostatted species and species that appear only as reactants or only as products, together with reactions that act as a source or sink. Thus, autonomous networks are natural candidates for the self-maintaining biochemical core of a cell, as they describe the internally closed reaction architecture required to sustain ongoing biochemical activity.
For autonomous networks, it can be shown that the MAF satisfies $0<\alpha<\infty$ (SM). A CRN is autocatalytic if, in addition to being autonomous, there exists a reaction flux such that all species are net produced (SM). From Equation 3, it follows immediately that $\alpha>1$ for autocatalytic networks. In SM, we show that the MAF of reducible networks is determined by their irreducible component with the highest MAF; thus, $\alpha$ acts as a “best-module” topological score. For a given network, efficient algorithms and their implementation for identifying irreducible autocatalytic subnetworks with the highest MAF can be found in [blanco2025identifyingselfamplifyinghypergraphstructures].
A network is considered to be minimal with respect to a certain property if it does not contain any subnetwork that also verify the property. In particular, finding the MAF for minimal autonomous networks reduces to solving the generalized-eigenvalue problem of the form $\mathbb{S}^{+}\mathbf{x}=\alpha\phantom{\cdot}\mathbb{S}^{-}\mathbf{x}.$ Minimal autocatalytic networks, termed (autocatalytic) cores, in reversible networks are shown to be of five distinct types [blokhuis2020universal], displayed in Fig. 1. These cores provide useful reference points for interpreting $\alpha$ as an amplification strength (see Table 1), yielding characteristic values such as $\sqrt{2}$ (Type I and III), the golden ratio $(\sqrt{5}-1)/2$ (Type IV) and $2$ (Type V). Numerical approximations of the exact values are summarized in Table 1. Since for minimal networks, Eq. 3 reduces to the generalized eigenvalue problem, the ‘dual’ autoinhibitory core resulting from reaction reversal exchanges $\mathbb{S}^{+}\leftrightarrow\mathbb{S}^{-}$ has MAF $1/\alpha$. Note that, however, for an arbitrary network, $\alpha(\mathbb{S}^{-},\mathbb{S}^{+})\neq 1/\alpha(\mathbb{S}^{+},\mathbb{S}^{-})$.
Figure 1: Five types of minimal autocatalytic cores (from Ref. [blokhuis2020universal]). Reversing all reactions exchanges $\mathbb{S}^{+}\leftrightarrow\mathbb{S}^{-}$ and maps the MAF as $\alpha\mapsto 1/\alpha$, yielding the corresponding dual autoinhibitory motif.
Type I Type II Type III Type IV Type V
$\alpha$ (core) 1.41 1.32 1.41 1.62 2.00
$\alpha$ (dual) 0.71 0.75 0.71 0.61 0.50
Table 1: Representative MAF values for autocatalytic cores and their autoinhibitory duals, consistent with [blokhuis2020universal].
Balanced growth in scalable reaction networks— We consider the regime of balanced growth, which is typically used to study chemical attractors [singh2023multistable] and metabolic networks [dourado_analytical_2020]. For species population vector $\mathbf{n}$, we introduce the extensive quantity $\Omega=\mathbf{w}^{T}\mathbf{n}$, where the vector $\mathbf{w}$ quantifies the relative weights of the species. Then, the relative population vector $\mathbf{q}=\mathbf{n}/\Omega$ is such that $\mathbf{w}^{T}\mathbf{q}=1.$ A choice of CRN kinetics $\mathbf{J}(\mathbf{y})$ is a Class-1 system [pandey2020exponential], or scalable [lin2020origin], if it satisfies $\mathbf{J}(\mathbf{y})/c=\mathbf{J}(\mathbf{y}/c),$ for any $c>0$. For scalable fluxes, the rescaled reaction flux $\mathbf{J}(\mathbf{n})/\Omega=\mathbf{J}(\mathbf{q})$ only depends on $\mathbf{q}$. Under this condition, the composition dynamics of the relative population on the simplex reads (see SM for details):
$\dot{\mathbf{q}}=\mathbb{S}\,\mathbf{J}(\mathbf{q})-\bigl(\mathbf{w}^{T}\mathbb{S}\,\mathbf{J}(\mathbf{q})\bigr)\mathbf{q}.$ (4)
Assuming that (4) admits a stable fixed point $\mathbf{q}^{*}$, the population scale evolves exponentially, $\dot{\Omega}=\Lambda\,\Omega$, with balanced-growth rate $\Lambda$ characterized by
$\mathbb{S}\,\mathbf{J}(\mathbf{q}^{*})=\Lambda\,\mathbf{q}^{*},\qquad\Lambda=\mathbf{w}^{T}\mathbb{S}\,\mathbf{J}(\mathbf{q}^{*}).$ (5)
We refer to a system with $\Lambda>0$, $\Lambda=0$, and $\Lambda<0$, respectively, as a ‘growing’, ‘steady’, and ‘shrinking’ system.
Figure 2: Growth parameter $\delta=\alpha-1$ versus normalized growth rate $\tilde{\Lambda}=\Lambda/\|\mathbb{S}^{-}\|_{k}$ for random autonomous, unambiguous CRNs with $2\leq|\mathcal{S}|\leq 12$ and $2\leq|\mathcal{R}|\leq 12$ and $M^{+}=M^{-}=2$. Left: autoinhibitory networks ($\alpha<1$, $\delta<0$). Right: autocatalytic networks ($\alpha>1$, $\delta>0$) and networks which admit a steady-state flux ($\alpha=1$, $\delta=0$). For $\alpha<1$, all points lie within the expected strip $1\leq\tilde{\Lambda}\leq 0$, and for $\alpha\geq 1$, the rigorous bound $1\leq\tilde{\Lambda}\leq\delta$ holds throughout. Furthermore, most autoinhibitory networks satisfy $\delta\leq\tilde{\Lambda}\leq 0$, while rare red outliers illustrate that the lower bound is not universal. The marginal distribution of $\delta$ is shown along the right side of each panel. The distribution exhibits accumulation near specific values, as discussed in the main text and Fig. 3.
Figure 3: Density of $\alpha$ over random networks from Fig. 2 with motif-labelled peaks. Left: autoinhibitory networks with $\alpha<1$. Right: autocatalytic networks with $\alpha>1$ and networks exhibiting a steady state with $\alpha=1$. Peaks are labeled by core types that share the same value of $\alpha$. For autocatalytic networks, some peaks correspond to autocatalytic cores, while other peaks correspond to non-minimal structures, such as the peak at $\alpha=1$ for non-autocatalytic structures or the peak at $\alpha=2$. For autoinhibitory networks, some peaks correspond to the reciprocal autoinhibitory cores, obtained by reversing the reaction arrows from their autocatalytic definitions (see Table 1).
Topological bounds on $\Lambda$ Consider scalable mass-action kinetics with reaction rates
$\mathbf{J}_{r}(\mathbf{n})$ $=k_{r}\frac{1}{\Omega^{\|(\mathbb{S}^{-})^{r}\|-1}}\prod_{s\in\mathcal{S}}\mathbf{n}_{s}^{(\mathbb{S}^{-})^{r}_{s}},$ (6)
where $k_{r}>0$ is the rate constant associated with reaction $r$, $\|(\mathbb{S}^{-})^{r}\|=\sum_{s}(\mathbb{S}^{-})^{r}_{s}$ is the total order of the reactants. In contrast to standard mass-action kinetics, where $\Omega$ denotes a constant volume, which has a simple expression for ideal gas-phase systems [gillespie1992rigorous], the choice $\Omega=\mathbf{w}^{T}\mathbf{n}$ is natural for modeling ideal fluids with $\mathbf{w}$ encoding the relative weights of the species contributing to the total volume.
We now focus on autonomous scalable mass-action reaction networks with $\mathbf{w}\geq\bm{1}$ that admit a stable fixed point $\mathbf{q}^{*}$ in the positive orthant such that $\mathbf{q}^{*}\in\mathbb{R}_{>0}^{|\mathcal{S}|}$. For such networks, $\mathbb{S}^{-}\mathbf{J}(\mathbf{q})\leq\|\mathbb{S}^{-}\|_{k}\phantom{\cdot}\mathbf{q},$ where the kinetic–stoichiometric prefactor is
$\|\mathbb{S}^{-}\|_{k}\;:=\;\max_{s\in\mathcal{S}}\;\sum_{r\in\mathcal{R}}(\mathbb{S}^{-})^{r}_{s}\,k_{r}.$ (7)
From this relation, we obtain a universal loose lower-bound $\|\mathbb{S}^{-}\|_{k}<\Lambda.$ We then show that if the MAF of a network is not greater than one, the network must either be steady or shrinking, i.e. $\alpha\leq 1\implies\Lambda\leq 0$. Finally, we show that growing networks satisfy $\Lambda\;\leq\;\|\mathbb{S}^{-}\|_{k}\,(\alpha-1).$ Collecting the above relations in a single equation, we get the main result of this work
$\|\mathbb{S}^{-}\|_{k}\;<\;\Lambda\;\leq\;\max\left(0,\|\mathbb{S}^{-}\|_{k}\,(\alpha-1)\right).$ (8)
The upper bound is saturated for autocatalytic networks with unimolecular inputs and equal rate constants. For these networks, furthermore, $(\alpha-1)$ coincides with the largest eigenvalue of the stoichiometric matrix $\mathbb{S}$. Full details of these results, along with a worked example, are provided in the SM.
Equation (8) turns the MAF $\alpha$ into a sharp, parameter-sparse descriptor of growth. First, it yields a strict decay certificate for non-autocatalytic networks: if $\alpha\leq 1$, then $\Lambda\leq 0$ independently of the rate constants and of the choice of stable fixed point. Second, it imposes an explicit upper bound for autocatalytic networks: even under arbitrary kinetic tuning, the balanced growth rate cannot exceed $\|\mathbb{S}^{-}\|_{k}(\alpha-1)$. While Eq. (8) implies that any growing network must be autocatalytic, i.e. $\Lambda>0\implies\alpha>1$, the converse does not hold: an autocatalytic network need not exhibit positive growth. In other words, the sign of $\Lambda$ fundamentally depends on the kinetics when $\alpha>1$.
When bounds on rate constants $k_{r}$ are available, Eq. (8) leads to a clear separation between stoichiometric and kinetic structure of the network. Indeed, the growth rate $\Lambda$, rescaled by $K=\max_{r}k_{r}$, is then bounded by quantities that depend only on the network topology:
$\|\mathbb{S}^{-}\|_{1}\;<\;\frac{\Lambda}{K}\;\leq\;\,\max\,\!\bigl(0,\|\mathbb{S}^{-}\|_{1}(\alpha-1)\bigr),$ (9)
where $\|\mathbb{S}^{-}\|_{1}=\max_{s}\sum_{r}(\mathbb{S}^{-})_{s}^{\,r}$.
Numerical tests on random autonomous networks— We tested the bounds on randomly generated autonomous, unambiguous CRNs with $2\leq|\mathcal{S}|\leq 12$ and $2\leq|\mathcal{R}|\leq 12$ (SM). We also vary the maximum reaction order for products, $M^{+}$, and reactants, $M^{-}$, where $M^{\pm}:=\max_{r\in\mathcal{R}}\sum_{s\in\mathcal{S}}(\mathbb{S}^{\pm})_{s}^{r}.$ For each instance, we computed $\alpha$ from (3) using computational implementations in [sargent2019neumann, blanco2025identifyingselfamplifyinghypergraphstructures]. When necessary, we also used integration of the scalable composition dynamics to its stable fixed point $\mathbf{q}^{*}$ following [lin2020origin], and extracted $\Lambda$ from (5). Using $\delta\equiv\alpha-1$ and $\tilde{\Lambda}\equiv\Lambda/\|\mathbb{S}^{-}\|_{k}$, Equation 8 can be recast as
$1<\tilde{\Lambda}\leq\max(0,\delta).$ (10)
Figure 2 confirms that all networks satisfy this bound. The majority of autocatalytic networks with ($\alpha>1$) are far from the limit $\delta\simeq 0$, and their growth rate stays far from the lower bound $\tilde{\Lambda}=-1$ and is positive. Thus, these networks have a built-in tendency for growth. The majority of autoinhibitory networks ($\alpha<1$), satisfy the empirical envelope $\delta\leq\tilde{\Lambda}$. This lower bound is much tighter than the one of Eq. (8), but it is not universal: it can fail. In that case, one obtains the outliers shown in red in the right figure of Fig. 2. As explained in SM, these outliers mainly appear in the region near $\delta\simeq 0^{-}$.
A striking feature of the left column of Fig. 2 for maximum reaction orders $M^{+}=M^{-}=2$ is that the distribution of $\delta=\alpha-1$ values concentrates on a discrete set of values. In Fig. 3, we have labeled the peaks in $\alpha$ together with a representation of their core structure. We observe that a large number of networks concentrate on the MAF value of type I or type III cores, which is $\sqrt{2}$ or on the value 2, which corresponds to the MAF value of type V. A large degeneracy of networks is observed among networks with $\alpha\geq 1$, at values $\alpha=1$, and at $\alpha=2$, which coincides with the value of $M^{+}$. In SM, we show that the MAF $\alpha$ satisfies $1/M^{-}\leq\alpha\leq M^{+}$. For higher reaction orders $M^{+}$ and $M^{-}$, the peaks flatten out and the distribution approaches a continuous spectrum.
Discussion and outlook— The bounds of Eq. (8) turn a purely structural quantity, the MAF $\alpha$, into a quantitative control parameter for balanced growth. Operationally, one can compute $\alpha$ from stoichiometry [blanco2025identifyingselfamplifyinghypergraphstructures, sargent2019neumann, de_martino_von_2012], then estimate $\|\mathbb{S}^{-}\|_{k}$ from input stoichiometries and rough kinetic scales, to obtain an upper bound on balanced growth.
An important observation is that, for a fixed topology, varying rate constants changes $\Lambda$ continuously, within limits set by $\alpha$. This suggests the following evolutionary scenario for the improvement in chemical growth. One route is kinetic tuning under fixed stoichiometric constraints, which moves $\Lambda$ within the admissible interval while the MAF $\alpha$ remains unchanged. The large degeneracy of networks observed for certain MAF values suggest the possibility of neutral evolution at fixed MAF value.
The second route is structural modification (adding/removing reactions, merging motifs, rewiring autocatalytic components), which changes $\alpha$ itself and thereby shifts the admissible bound. In evolutionary terms, this means a punctuated equilibrium with topological network changes playing the role of mutations that alter the admissible growth bound, while subsequent kinetic adaptation and neutral evolution adjust $\Lambda$ within the available range [qian_chemical_2010, sole_fundamental_2024, jain_model_2001].
Beyond this scenario, our result has also practical implications for open chemical reactors because in a continuously stirred reactor with dilution rate $d$, the population scale following a possible transient satisfies $\dot{\Omega}=(\Lambda-d)\,\Omega$. Then, sustained growth requires $\Lambda>d$, therefore $\Lambda$ acts as a dilution threshold above which the system is washed out. Combined with Eq. (8), this yields a necessary structural condition for persistence against dilution: even with optimal kinetic tuning, survival at dilution $d$ is excluded unless $\alpha>1$ and $d$ lies below the corresponding growth bound [kolchinsky2025thermodynamics].
In compartmentalized systems with a steady-composition, the growth rate $\Lambda$ controls the rate of change of any extensive quantity besides that of the global population [sughiyama2022chemical]. The compartment volume is one such quantity, and indeed, experiments have shown that the compartment volume can inform on the autocatalytic growth of small confined chemical species and that it represents a proxy for fitness when coupling many such compartments together [lu_small-molecule_2024].
These interpretations suggest several natural extensions of the present work. First, it opens a route towards network engineering: using only knowledge of stoichiometry, one could design networks, that could potentially drive growth above a target threshold, or induce a robust decay for $\alpha\leq 1$ (robust decay). Pursuing this program requires composition rules for $\alpha$: which elementary network operations (adding a reaction, merging species, etc.) increase or decrease $\alpha$ [sarkar_design_2019, despons_stability_2025], and by how much? For related results in the economics literature, see [bapat1995perron]. Such a set of design rules would make $\alpha$ predictive under network modification and enable sharp statements about minimal structural changes needed to shift admissible growth bounds.
A second direction is to close the gap between topological capability ($\alpha$) and dynamical realization (the steady flux). Identifying sufficient dynamical conditions that force the steady flux to approach MAF-optimal patterns would turn the universal bound into a quantitative estimate for restricted kinetic classes and clarify which motifs control performance. At the same time, an improved understanding of the connection between the stoichiometric description of CRNs and thermodynamics [despons_structural_2024] could uncover functional organization principles in CRNs energetics. While $\alpha$ is purely stoichiometric, sustained growth in open chemistry requires chemical work supplied by chemostats [liang2024thermodynamic, gagrani2025thermodynamic, marehalli_srinivas_thermodynamics_2024]. A natural extension is therefore to connect growth bounds to thermodynamic budgets stated in terms of chemostat chemical potentials, affinities, or dissipation rates.
Finally, scalable balanced growth is an asymptotic statement. In finite populations, demographic noise and rare-event switching can induce extinctions or transitions between composition attractors even when a stable $\mathbf{q}^{*}$ exists deterministically [singh2023multistable, lazarescu_large_2019]. Extending the framework to stochastic scalable kinetics would quantify the reliability of a topologically permitted growth bounds and provide risk-aware structural criteria, such as expected persistence times under fluctuations.
Acknowledgments— P.G. acknowledges Barnabé Ledoux, Eric Smith, Nathaniel Virgo, and Atsushi Kamimura for discussions. P.G. and D.L. acknowledge P.G. was funded by JST CREST (JPMJCR2011) and JSPS Grant No. 25H01365. J.W. acknowledges Chunqiu Xia for computational resources. This research was partially conducted while visiting the Okinawa Institute of Science and Technology (OIST) through the Theoretical Sciences Visiting Program (TSVP) Thematic Program on Biological Information Processing (TP24BI).
Supplemental Material for
“Topological bounds on the dynamical growth rate of chemical reaction networks”
Supplemental Material for “Topological bounds on the dynamical growth rate of chemical reaction networks”
A. Structural setting and definitions
A1. Notation
Vectors are represented in boldface. The $i$th component of a vector $\mathbf{w}$ is denoted with $\mathbf{w}_{i}$. A vector $\mathbf{w}$ is said to be greater than zero, denoted as $\mathbf{w}>\mathbf{0}$, if $\mathbf{w}_{i}\geq 0$ for all $i$ but $\mathbf{w}\neq\mathbf{0}$. A vector $\mathbf{w}$ is said to be strictly greater than zero, denoted as $\mathbf{w}\gg\mathbf{0}$, if $\mathbf{w}_{i}>0$ for all $i$. Each vector is assumed to be a column vector. Matrices are denoted in blackboard bold. The entry of the $i$th column and $j$th row of the matrix $\mathbb{A}$ is denoted as $\mathbb{A}^{i}_{j}$. A matrix $\mathbb{A}$ is said to be nonnegative if all its entries are nonnegative.
A2. Chemical reaction network
A chemical reaction network (CRN) with species set $\mathcal{S}=\{1,\dots,|\mathcal{S}|\}$ and reaction set $\mathcal{R}=\{1,\dots,|\mathcal{R}|\}$ is encoded by two nonnegative matrices $|\mathcal{S}|\times|\mathcal{R}|$ matrices
$(\mathbb{S}^{-})_{s}^{r}\geq 0,\qquad(\mathbb{S}^{+})_{s}^{r}\geq 0,$ (S1)
collecting the input (reactant) and output (product) stoichiometric coefficients of reaction $r$ on species $s$. In this work, we restrict the stoichiometries to be integers. The stoichiometric matrix is
$\mathbb{S}=\mathbb{S}^{+}-\mathbb{S}^{-}.$ (S2)
For any nonnegative flux vector $\mathbf{x}\geq\mathbf{0}\in\mathbb{R}^{|\mathcal{R}|}_{\geq 0}$, the species-wise input and output fluxes are
$\mathbb{S}^{-}\mathbf{x}\geq\mathbf{0}\in\mathbb{R}^{|\mathcal{S}|}_{\geq 0},\qquad\mathbb{S}^{+}\mathbf{x}\geq\mathbf{0}\in\mathbb{R}^{|\mathcal{S}|}_{\geq 0}$. (S3)
A3. Autonomous network
A CRN $(\mathbb{S}^{+},\mathbb{S}^{-})$ is autonomous if [gagrani2024polyhedral, blanco2025identifyingselfamplifyinghypergraphstructures]:
  1. (A1)
    All species are consumed by some reaction.
    $\forall s\in\mathcal{S},\exists r\in\mathcal{R}:(\mathbb{S}^{-})^{r}_{s}>0.$
  2. (A2)
    All species are produced by some reaction.
    $\forall s\in\mathcal{S},\exists r\in\mathcal{R}:(\mathbb{S}^{+})^{r}_{s}>0.$
  3. (A3)
    All reactions consume some species.
    $\forall r\in\mathcal{R},\exists s\in\mathcal{S}:(\mathbb{S}^{-})^{r}_{s}>0.$
  4. (A4)
    All reactions produce some species.
    $\forall r\in\mathcal{R},\exists s\in\mathcal{S}:(\mathbb{S}^{+})^{r}_{s}>0.$
A network where each reaction consumes and produces each species, i.e. which satisfies conditions (A2) and (A3), will be called semi-autonomous.
A4. Autocatalytic network
A CRN $(\mathbb{S}^{+},\mathbb{S}^{-})$ is autocatalytic if:
  1. (AC1)
    Autonomy: The CRN is autonomous (satisfies (A1)-(A4)).
  2. (AC2)
    Productivity: There exists a flux vector $\mathbf{x}$ such that all species are net-produced positively,
    $\exists\mathbf{x}>\mathbf{0}:\;\mathbb{S}\mathbf{x}\gg\mathbf{0}\iff\mathbb{S}^{+}\mathbf{x}\gg\mathbb{S}^{-}\mathbf{x}\iff\alpha(\mathbf{x})>1.$ (S4)
Note that requiring all reactions to be unambiguous-i.e. no species are shared between the reactants and products of any reaction-recovers the corresponding definitions of autonomy and autocatalysis in [blokhuis2020universal]. The results in this work does not require this assumption.
A5. Reducible and irreducible network
A CRN is reducible [gale1989theory, sargent2019neumann] if there exists a semi-autonomous subnetwork inside the whole network, that is able to produce every species in itself without consuming species from outside. Mathematically, a CRN $(\mathcal{S},\mathcal{R})$ is reducible if there exists $(\mathcal{S}^{\prime},\mathcal{R}^{\prime})$ such that
$\forall s\in\mathcal{S}^{\prime},\exists r\in\mathcal{R}^{\prime}:$ $(\mathbb{S}^{+})^{r}_{s}>0,$ (S5)
$\forall r\in\mathcal{R}^{\prime},\forall s\in\mathcal{S}/\mathcal{S}^{\prime}:$ $(\mathbb{S}^{-})^{r}_{s}=0.$ (S6)
A CRN is said to be irreducible if it is not reducible.
B. Maximum Amplification Factor (MAF) results
S0.1 B1. Reformulation of MAF
The MAF can be reformulated as the solution to the following optimization problem:
$\alpha(\mathbb{S}^{+},\mathbb{S}^{-})=$ $\hskip-50.00008pt\max\limits_{\mathbf{x}>\mathbf{0}}\qquad\alpha(\mathbf{x})$ (S7)
$\text{such that }\quad\mathbb{S}^{+}\mathbf{x}\geq\alpha(\mathbf{x})\phantom{\cdot}\mathbb{S}^{-}\mathbf{x}.$
For solver implementations of the above, see [sargent2019neumann, blanco2025identifyingselfamplifyinghypergraphstructures].
B2. MAF bounds for autonomous networks
Von Neumann in [neumann1945model] and Theorem 9.8 in [gale1989theory] rigorously prove that $0<\alpha<\infty$ for semi-autonomous networks. Since autonomous networks are also semi-autonomous, the bounds trivially extend to them as well. While detailed proofs of the statement can be found there, we sketch a heuristic derivation of the same.
Recall from Equation 3 that
$\alpha\;=\;\max_{\mathbf{x}>\bm{0}}\;\min_{s\in\mathcal{S}}\frac{(\mathbb{S}^{+}\mathbf{x})_{s}}{(\mathbb{S}^{-}\mathbf{x})_{s}}.$ (S8)
Since, by property (A2), every species is produced by at least one reaction, there exists a feasible flux $\mathbf{x}>\mathbf{0}$ such that the numerator $\mathbb{S}^{+}\mathbf{x}\gg\mathbf{0}$. Thus, for such a flux, $(\mathbb{S}^{+}\mathbf{x})_{s}/(\mathbb{S}^{-}\mathbf{x})_{s}>0$ for every $s$ and, in particular, the minimum over all species is greater than zero. Furthermore, by property (A3), every reaction consumes and thus, for the same $\mathbf{x}$, $\mathbb{S}^{-}\mathbf{x}>\mathbf{0}$ (as not all elements $(\mathbb{S}^{-}\mathbf{x})_{s}$ can be simultaneously zero). Thus, there must exist at least one species $s$ for which $(\mathbb{S}^{+}\mathbf{x})_{s}/(\mathbb{S}^{-}\mathbf{x})_{s}<\infty$ which implies that minimum over all species must also be less than infinity. Since $\alpha$ is the maximum over such values, it must also satisfy $0<\alpha<\infty$.
B3. MAF of reducible networks
Von Neumann in [neumann1945model] and Theorem 9.10 in [gale1989theory] prove that there exists a unique MAF for irreducible semi-autonomous networks. Since autonomous networks are also semi-autonomous, irreducible autonomous networks have a unique MAF.
We now turn to the case of reducible networks. If the network decomposes into disjoint irreducible components that do not exchange mass—equivalently, if $(\mathbb{S}^{+},\mathbb{S}^{-})$ can be brought into block-triangular form after a suitable permutation—then the maximum amplification factor is simply the maximum of the MAFs of the individual irreducible blocks.
More generally, a reducible network may consist of an irreducible core whose species feed one or more downstream subnetworks. In this case, any flux through downstream reactions necessarily increases the consumption of species in the core without providing compensating production within it. As a result, activating downstream subnetworks cannot increase the uniform production-to-consumption ratio across species, and can only reduce the achievable amplification factor. Optimal fluxes therefore concentrate on the dominant irreducible component.
In this sense, $\alpha$ functions as a “best-module” structural score and is determined solely by the irreducible component with the highest MAF.
B4. MAF of minimal networks
Observe from the second formulation of the MAF (Equation S7) that the optimal solution $\mathbf{x}>\mathbf{0}$ for a CRN $(\mathbb{S}^{+},\mathbb{S}^{-})$ satisfies
$\mathbb{S}^{+}\mathbf{x}\;\geq\;\alpha\,\mathbb{S}^{-}\mathbf{x}.$ (S9)
While, in general, the above equation only holds as an inequality, the equality must be satisfied for some species and reactions otherwise $\alpha$ would not be the maximum value that satisfies the inequality. Given the subnetwork $(\underline{\mathbb{S}}^{+},\underline{\mathbb{S}}^{-})$ and $\underline{\mathbf{x}}$, we obtain the following generalized eigenvalue problem:
$\underline{\mathbb{S}}^{+}\underline{\mathbf{x}}=\alpha\,\underline{\mathbb{S}}^{-}\underline{\mathbf{x}},$ (S10)
with the constraint $\underline{\mathbf{x}}>\mathbf{0}$. The relation between solutions of Eq. (S7) and the generalized eigenvalue problem is studied in Ref. [thompson1971neumann].
An autonomous network is minimal if there is no subnetwork that is also autonomous. Since there is no species that is never consumed, such a network must also be irreducible. It has been shown that minimal autonomous networks have an equal number of species and reactions [gagrani2024polyhedral, blokhuis2020universal, vassena2024unstable], and that their input matrix can be rearranged into a diagonal form. Clearly, then, the network is irreducible and $\mathbb{S}^{-}$ is non-singular. Using the results in [thompson1971neumann], the problem of finding MAF reduces to the generalized eigenvalue problem of Eq. (S10), which further reduces to the standard eigenvalue problem
$(\mathbb{S}^{-})^{-1}\mathbb{S}^{+}\mathbf{x}=\alpha\,\mathbf{x},$ (S11)
where $\alpha$ is the highest eigenvalue of $(\mathbb{S}^{-})^{-1}\mathbb{S}^{+}$ with a corresponding nonnegative Perron-Frobenius eigenvector $\mathbf{x}$.
C. Composition dynamics
For species population $\mathbf{n}$, relative weight vector $\mathbf{w}>\mathbf{0}$, and scale $\Omega=\mathbf{w}^{T}\mathbf{n}$, we define the relative population vector
$\mathbf{q}$ $=\frac{\mathbf{n}}{\Omega}=\frac{\mathbf{n}}{\mathbf{w}^{T}\mathbf{n}}.$ (S12)
Consider a scalable dynamical system
$\dot{\mathbf{n}}(t)$ $=\mathbb{S}\mathbf{J}(\mathbf{n}(t)),$ (S13)
that satisfies
$\frac{\mathbf{J}(\mathbf{n})}{\Omega}=\frac{\mathbf{J}(\mathbf{q}\Omega)}{\Omega}$ $=\mathbf{J}(\mathbf{q}),$ (S14)
Taking a time derivative of $\Omega$ and using the change of coordinates readily yields
$\dot{\Omega}=\Omega\,\mathbf{w}^{T}\mathbb{S}\mathbf{J}(\mathbf{q}).$ (S15)
Substituting $\mathbf{n}=\mathbf{q}\Omega$ and using the chain-rule in Equation S13 yields the composition equation
$\dot{\mathbf{q}}$ $=\mathbb{S}\mathbf{J}(\mathbf{q})-(\mathbf{w}^{T}\mathbb{S}\mathbf{J}(\mathbf{q}))\mathbf{q}.$ (S16)
The fixed point of the above equation yields the relation stipulated in Equation 5 of the main text.
D. Bounds on the dynamic growth rate
D1. Assumptions
We assume an autonomous network under mass–action kinetics Eq. (6) with positive rate constants and the existence of at least one stable fixed point of Equation S16 in the positive orthant $\mathbf{q}^{*}\gg\mathbf{0}$, ensuring that the growth is exponential with a rate defined by Eq. (5).
We also assume that the relative weight vector satisfy $\mathbf{w}\geq\mathbf{1}$, which simply means that the population of each species has a positive contribution in $\Omega$. Since $\mathbf{w}^{T}\mathbf{q}=1$, it follows that $\mathbf{q}\leq 1$.
S0.2 D2. Kinetic-stoichiometric prefactor
From the form of the propensity of Eq. (6),
$(\mathbb{S}^{-}\mathbf{J}(\mathbf{q}))_{s}:=\sum_{r\in\mathcal{R}}(\mathbb{S}^{-})^{r}_{s}k_{r}\prod_{t\in\mathcal{S}}\mathbf{q}_{t}^{(\mathbb{S}^{-})^{r}_{t}}.$ (S17)
Given assumption (A1), for each species $s$, there exists a reaction $r$ such that $(\mathbb{S}^{-})^{r}_{s}\geq 1$. For such a pair of $s$ and $r$, we rewrite $\mathbf{q}^{(\mathbb{S}^{-})^{r}_{s}}=\mathbf{q}_{s}p_{r}^{s}(\mathbf{q})$, where $p_{r}^{s}(\mathbf{q})$ is a monomial of the concentration coordinates. This yields,
$(\mathbb{S}^{-}\mathbf{J}(\mathbf{q}))_{s}$ $=\mathbf{q}_{s}\sum_{r:(\mathbb{S}^{-})_{s}^{r}>0}(\mathbb{S}^{-})^{r}_{s}k_{r}p^{s}_{r}(\mathbf{q}).$ (S18)
Since all the concentrations are upper bounded by 1, for every $r$, $p^{s}_{r}(\mathbf{q})\leq 1$. Putting these together, we get
$(\mathbb{S}^{-}\mathbf{J}(\mathbf{q}))_{s}\leq\mathbf{q}_{s}\sum_{r}(\mathbb{S}^{-})^{r}_{s}k_{r}\leq\mathbf{q}_{s}\max_{t\in\mathcal{S}}\left(\sum_{r}(\mathbb{S}^{-})^{r}_{t}k_{r}\right).$ (S19)
Using
$\|\mathbb{S}^{-}\|_{k}\;:=\;\max_{s\in\mathcal{S}}\;\sum_{r\in\mathcal{R}}(\mathbb{S}^{-})^{r}_{s}\,k_{r},$ (S20)
we obtain the inequality
$\mathbb{S}^{-}\mathbf{J}(\mathbf{q})\leq\|\mathbb{S}^{-}\|_{k}\phantom{\cdot}\mathbf{q}.$ (S21)
S0.3 D3. Universal lower bound
By Equation 5, $\mathbb{S}\mathbf{J}(\mathbf{q}^{*})=\Lambda\mathbf{q}^{*}$, for every species $s$
$\Lambda=\frac{(\mathbb{S}\mathbf{J}(\mathbf{q}^{*}))_{s}}{\mathbf{q}^{*}_{s}}=\frac{(\mathbb{S}^{+}\mathbf{J}(\mathbf{q}^{*}))_{s}-(\mathbb{S}^{-}\mathbf{J}(\mathbf{q}^{*}))_{s}}{\mathbf{q}^{*}_{s}}.$ (S22)
Since $\mathbf{q}^{*}$ is in the positive-orthant, $\mathbf{J}(\mathbf{q}^{*})\gg\mathbf{0}$. Furthermore, by assumptions (A2) and (A4), every reaction produces and every species is produced by some reaction, yielding $\mathbb{S}^{+}\mathbf{J}(\mathbf{q}^{*})\gg\mathbf{0}$. This means that
$\Lambda>-\frac{(\mathbb{S}^{-}\mathbf{J}(\mathbf{q}^{*}))_{s}}{\mathbf{q}^{*}_{s}}\geq-\|\mathbb{S}^{-}\|_{k},$ (S23)
where Eq. (S21) is used in the last line.
D4. $\alpha\leq 1$ implies $\Lambda\leq 0$
From the definitions of MAF, Equations 2 and 3 in the main text, we obtain the following implication and its contrapositive:
$\exists\mathbf{x}>\mathbf{0}:$ $\;\mathbb{S}\mathbf{x}\gg\mathbf{0}\implies 1<\alpha(\mathbf{x})\leq\alpha,$
$\nexists\mathbf{x}>\mathbf{0}:$ $\;\mathbb{S}\mathbf{x}\gg\mathbf{0}\impliedby\alpha(\mathbf{x})\leq\alpha\leq 1.$ (S24)
Since $\Lambda$ is of the form $\mathbf{w}^{T}\mathbb{S}\mathbf{x}$, we can see from the second line in Equation S24 that $\alpha\leq 1\implies\Lambda\leq 0$, i.e. if the MAF of a network is not greater than one, the network must either be steady or shrinking.
D5. Upper bound on growing networks
For growing networks, for which $\Lambda>0$,
$\Lambda$ $=\min_{s\in\mathcal{S}}\frac{(\mathbb{S}\mathbf{J}(\mathbf{q}^{*}))_{s}}{(\mathbb{S}^{-}\mathbf{J}(\mathbf{q}^{*}))_{s}}\frac{(\mathbb{S}^{-}\mathbf{J}(\mathbf{q}^{*}))_{s}}{\mathbf{q}^{*}_{s}}$
$\leq\|\mathbb{S}^{-}\|_{k}\min_{s\in\mathcal{S}}\frac{\left(\mathbb{S}\mathbf{J}(\mathbf{q}^{*})\right)_{s}}{\left(\mathbb{S}^{-}\mathbf{J}(\mathbf{q}^{*})\right)_{s}},$
$\leq\|\mathbb{S}^{-}\|_{k}\max_{\mathbf{x}>\mathbf{0}}\min_{s\in\mathcal{S}}\frac{(\mathbb{S}\mathbf{x})_{s}}{(\mathbb{S}^{-}\mathbf{x})_{s}}$
$\leq\|\mathbb{S}^{-}\|_{k}\left(\max_{\mathbf{x}>\mathbf{0}}\min_{s\in\mathcal{S}}\frac{(\mathbb{S}^{+}\mathbf{x})_{s}}{(\mathbb{S}^{-}\mathbf{x})_{s}}-1\right)$
$\leq\|\mathbb{S}^{-}\|_{k}\left(\alpha(\mathbb{S}^{+},\mathbb{S}^{-})-1\right),$ (S25)
where in the second line $\Lambda>0$ and Equation S21 are used, and the definition of MAF (Eq. 3) is used to obtain the final line. Observe that, by assumptions (A1) and (A3), every reaction consumes and every species is consumed by some reaction, yielding $\mathbb{S}^{-}\mathbf{J}(\mathbf{q}^{*})\gg\mathbf{0}$ and a positive denominator for every $s$.
D6. Non-universal lower bound
For shrinking networks, for which $\Lambda<0$, then:
$(\mathbb{S}^{+}\jmath(\mathbf{q}^{*}))_{s}<(\mathbb{S}^{-}\jmath(\mathbf{q}^{*}))_{s}.$ (S26)
We also have that $\Lambda$ can be expressed as:
$\Lambda=\min_{s\in\mathcal{S}}\frac{(\mathbb{S}\jmath(\mathbf{q}^{*}))_{s}}{q_{s}^{*}}=\min_{s\in\mathcal{S}}\left(\frac{(\mathbb{S}^{+}\jmath(\mathbf{q}^{*}))_{s}}{q_{s}^{*}}-\frac{(\mathbb{S}^{-}\jmath(\mathbf{q}^{*}))_{s}}{q_{s}^{*}}\right).$ (S27)
Using the definition of the MAF $\alpha$, for the flux vector $\jmath(\mathbf{q}^{*})$, we have:
$\alpha\geq\min_{s\in\mathcal{S}}\frac{(\mathbb{S}^{+}\jmath(\mathbf{q}^{*}))_{s}}{(\mathbb{S}^{-}\jmath(\mathbf{q}^{*}))_{s}}=\alpha(\jmath(\mathbf{q}^{*})).$ (S28)
Defining a positive quantity
$\epsilon:=\alpha-\alpha(\jmath(\mathbf{q}^{*})),$ (S29)
for each species $s$, we can write:
$\frac{(\mathbb{S}^{+}\jmath(\mathbf{q}^{*}))_{s}}{(\mathbb{S}^{-}\jmath(\mathbf{q}^{*}))_{s}}\geq\alpha-\epsilon.$ (S30)
Note that $\epsilon$ tends to zero as $\jmath(\mathbf{q}^{*})$ as approaches the optimal flux vector.
Using the above inequality, we can write:
$\Lambda\geq\min_{s\in\mathcal{S}}\left((\alpha-\epsilon-1)\frac{(\mathbb{S}^{-}\jmath(\mathbf{q}^{*}))_{s}}{q_{s}^{*}}\right).$ (S31)
Now using the inequality of Eq. (S21), namely $(\mathbb{S}^{-}\jmath(\mathbf{q}^{*}))_{s}/q_{s}^{*}\leq\|\mathbb{S}^{-}\|_{k}$, we have:
$\Lambda\geq(\alpha-\epsilon-1)\|\mathbb{S}^{-}\|_{k}.$ (S32)
For a majority of random networks, we empirically find
$\Lambda\geq(\alpha-1)\|\mathbb{S}^{-}\|_{k}.$ (S33)
This holds when $\jmath(\mathbf{q}^{*})$ is sufficiently close to the optimal flux vector, so that $\epsilon$ is negligible. However, there are some networks, especially those for which $\alpha-1$ is small, for which the above bound is broken. In other words, this lower bound is not universal. The networks that break the bound correspond to the outliers visible in Fig. 2.
E. Examples and numerical tests
S0.4 E1. Unimolecular inputs
From the discussion in Section B3, for minimal autonomous networks with unimolecular inputs, the input matrix is the identity, $\mathbb{S}^{-}=\mathbb{I}$, and the generalized eigenvalue problem reduces to $\mathbb{S}^{+}\mathbf{x}=\alpha\mathbf{x},$ where $\alpha$ is the Perron–Frobenius eigenvalue of $\mathbb{S}^{+}$. Furthermore, the same eigenvector $\mathbf{x}$ satisfies
$\mathbb{S}\mathbf{x}=(\mathbb{S}^{+}-\mathbb{I})\mathbf{x}=(\alpha-1)\mathbf{x},$ (S34)
from which we see that $(\alpha-1)$ is the largest eigenvalue of the stoichiometric matrix $\mathbb{S}$.
Moreover, from Equation 7,
$\|\mathbb{S}^{-}\|_{k}=\max_{r}k_{r},$ (S35)
so the kinetic–stoichiometric prefactor reduces to the maximum rate constant. Thus, in the case where all rate constants are equal, $k_{r}=k\ \forall r\in\mathcal{R}$, the inequality becomes an equality,
$\Lambda=k(\alpha-1),$ (S36)
and the fixed point $\mathbf{q}^{*}$ is given by the Perron–Frobenius eigenvector $\mathbf{x}$ of $\mathbb{S}$, normalized such that $\mathbf{w}^{T}\mathbf{x}=1$.
S0.5 E2. Worked example
Figure S1: The dynamical growth rate $\Lambda$ is shown for different choices of the rate constants, together with an overlay of the quantity $\|\mathbb{S}^{-}\|_{k}\bigl(\alpha(\mathbb{S}^{+},\mathbb{S}^{-})-1\bigr)$ for the type 1 autocatalytic core. In all cases, this quantity provides an upper bound on the growth rate $\Lambda$, and the bound is saturated when $k_{1}=k_{2}=1$.
Consider the CRN
$r_{1}:\text{A}\to\text{B},\qquad r_{2}:\text{B}\to 2\text{A}.$ (S37)
Then (ordering species as (A,B) and reactions as ($r_{1},r_{2}$)),
$\mathbb{S}^{+}$ $=\begin{bmatrix}0&2\\ 1&0\end{bmatrix}$, $\mathbb{S}^{-}$ $=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$, $\mathbb{S}$ $=\begin{bmatrix}-1&2\\ 1&-1\end{bmatrix}$. (S38)
Following Eq. 3,
$\alpha(\mathbb{S}^{+},\mathbb{S}^{-})=\max_{\mathbf{x}>\mathbf{0}}\min\left(\frac{2\mathbf{x}_{2}}{\mathbf{x}_{1}},\frac{\mathbf{x}_{1}}{\mathbf{x}_{2}}\right).$ (S39)
For any $\mathbf{x}_{2}$, it is clear that the max-min is attained at $\mathbf{x}_{1}=\sqrt{2}\mathbf{x}_{2}$. Thus,
$\alpha(\mathbb{S}^{+},\mathbb{S}^{-})$ $=\sqrt{2},$ $\mathbf{x}_{\alpha}=[\sqrt{2},1]^{T}.$ (S40)
Observe that $\alpha=\sqrt{2}$ is simply the highest eigenvalue of $\mathbb{S}^{+}$, and $(\sqrt{2}-1)$ is the highest eigenvalue of $\mathbb{S}$.
Under mass-action with $\mathbf{w}=[1,1]^{T}$ (so $\mathbf{q}_{A}+\mathbf{q}_{B}=1$) and $k_{1}=k_{2}=1$, the growth rate solution satisfies
$\Lambda$ $=\sqrt{2}-1,$ $\mathbf{q}_{A}^{*}$ $=\frac{\sqrt{2}}{\sqrt{2}+1},$ $\mathbf{q}_{B}^{*}$ $=\frac{1}{\sqrt{2}+1}.$ (S41)
For this assignment $\|\mathbb{S}^{-}\|_{k}=\max(1,1)=1$, and from Equation 8, we have the upper bound
$\Lambda\leq 1\cdot(\sqrt{2}-1),$ (S42)
which is exactly saturated. The dynamical growth rate overlaid with the theoretical upper bound is shown in Fig. S1.
S0.6 E3. Numerical construction of the random networks
We construct a random stoichiometric matrix $\mathbb{S}\in\mathbb{Z}^{S\times R}$, where $S$ and $R$ respectively denote the number of species and reactions in the network. The integers $S$ and $R$ are randomly drawn in the set $[2,12]$, where $12$ is chosen arbitrarily to restrict the maximum system size.
The stoichiometric matrix is decomposed in the output and input matrices $\mathbb{S}=\mathbb{S}^{+}-\mathbb{S}^{-}$. We define the product and reactant reaction order, respectively, as
$m_{r}^{+}:=\sum_{s\in\mathcal{S}}(\mathbb{S}^{+})^{r}_{s},\quad m_{r}^{-}:=\sum_{s\in\mathcal{S}}(\mathbb{S}^{-})^{r}_{s}.$ (S43)
Let us also introduce a couple of integer parameters $(M^{+},M^{-})$ referred to as the maximum forward (resp. backward) reaction order such that
$m_{r}^{+}\leq M^{+},\qquad m_{r}^{-}\leq M^{-}\quad\text{for all }r\in\mathcal{R}.$ (S44)
For each forward reaction $r$, the product reaction order $m_{r}^{+}$ is sampled uniformly in the set $1,M^{+}]$. Conditional on this value, we construct the column $(\mathbb{S}^{+})^{r}$ by repeatedly sampling a species uniformly until the first constraint of Eq. (S43) is satisfied. We also make sure that the maxima $(M^{+},M^{-})$ are reached by some reaction. This procedure yields the output matrix $\mathbb{S}^{+}$. Analogously, we also obtain the input matrix $\mathbb{S}^{-}$. The rate constants for a realization are picked from a uniform distribution between 0 and 1, and normalized by scaling the largest rate constant to 1.
Finally, only networks satisfying the unambiguous and autonomous conditions are retained. A network is unambiguous if no reaction shares a species between its reactant and product complexes. Although the theoretical results of this work do not require unambiguity, ambiguous reactions are autonomous in the repeated species. Such reactions therefore contain an irreducible autonomous subnetwork supported entirely on that species, which can generate trivial or structurally decoupled contributions to the MAF. By excluding these cases, we focus on networks in which amplification arises from genuine inter-species structure rather than from embedded single-species self-coupling motifs.
E4. Bounds on the MAF
By definition of the reaction order, we have
$\forall\,r:\;\sum_{s}(\mathbb{S}^{\pm})_{s}^{r}\leq M^{\pm}.$ (S45)
By the definition of the reaction order and the property of autonomy (every reaction consumes and produces at least one species), we obtain the right and left inequalities, respectively, in the following equation:
$\forall\,r:\;1\leq\sum_{s}(\mathbb{S}^{\pm})_{s}^{r}\leq M^{\pm}.$ (S46)
Consider the quantity
$Q(\mathbf{x}):=\frac{\sum_{s,r}(\mathbb{S}^{+})_{s}^{r}\mathbf{x}_{r}}{\sum_{s,r}(\mathbb{S}^{-})_{s}^{r}\mathbf{x}_{r}}.$ (S47)
By Eq. S46, we have
$\sum_{r}\mathbf{x}_{r}\leq\sum_{s,r}(\mathbb{S}^{\pm})_{s}^{r}\mathbf{x}_{r}\leq M^{\pm}\sum_{r}\mathbf{x}_{r},$ (S48)
which implies
$\frac{1}{M^{-}}\leq Q(\mathbf{x})\leq M^{+}.$ (S49)
Recall that the MAF $\alpha$ is defined as the maximum over $\mathbf{x}>\mathbf{0}$ of the minimum ratios $(\mathbb{S}^{+}\mathbf{x})_{s}/(\mathbb{S}^{-}\mathbf{x})_{s}$. Suppose $\alpha>M^{+}$, and there exists $\mathbf{x}$ such that $\forall\,s:\,(\mathbb{S}^{+}\mathbf{x})_{s}/(\mathbb{S}^{-}\mathbf{x})_{s}>M^{+}.$ Clearly,
$Q(\mathbf{x})=\frac{\sum_{s}(\mathbb{S}^{+}\mathbf{x})_{s}}{\sum_{s}(\mathbb{S}^{-}\mathbf{x})_{s}}>M^{+},$ (S50)
which would violate the right-hand side of Eq. S49 and yield a contradiction. Thus $\alpha\leq M^{+}$.
Now suppose $\alpha<1/M^{-}$ and there exists $\mathbf{x}^{*}$ such that
$\min_{s}\frac{(\mathbb{S}^{+}\mathbf{x}^{*})_{s}}{(\mathbb{S}^{-}\mathbf{x}^{*})_{s}}<\frac{1}{M^{-}}.$ (S51)
Following the earlier discussion around Eq. S10 of the solution for $\alpha$, there is a set of species $\mathcal{S}^{\prime}$ with $|\mathcal{S}^{\prime}|\geq 1$ in an irreducible subnetwork of the CRN such that the minimum is attained simultaneously. Defining
$Q^{\prime}(\mathbf{x}):=\frac{\sum_{s\in\mathcal{S}^{\prime},r}(\mathbb{S}^{+})_{s}^{r}\mathbf{x}_{r}}{\sum_{s\in\mathcal{S}^{\prime},r}(\mathbb{S}^{-})_{s}^{r}\mathbf{x}_{r}},$ (S52)
and using the same arguments as before, we see that $\forall\,\mathbf{x}:\,1/M^{-}\leq Q^{\prime}(\mathbf{x})$. However, if $\alpha<1/M^{-}$, then $Q^{\prime}(\mathbf{x}^{*})<1/M^{-}$, which violates the inequality for $Q^{\prime}(\mathbf{x})$. Again we obtain a contradiction, and thus $\alpha\geq 1/M^{-}$.
Putting all the previously stated results together, we obtain
$\frac{1}{M^{-}}\leq\alpha\leq M^{+}.$ (S53)
In what follows, we present examples of networks that saturate these bounds. For ambiguous networks with $|\mathcal{S}|=1$, the lower and upper bounds are saturated by the following networks, respectively:
$\{(M^{-})X\to X\}\quad\text{and}\quad\{X\to(M^{+})X\}.$ (S54)
An unambiguous network with $M^{+}=2$ and $|\mathcal{S}|=2$ that saturates the upper bound is (the construction for the lower bound is analogous):
$X_{1}$ $\to 2X_{2},$ (S55)
$X_{2}$ $\to 2X_{1}.$ (S56)
Similarly, an unambiguous network with $M^{+}=2$ and $|\mathcal{S}|=3$ that saturates the upper bound is:
$X_{1}$ $\to X_{2}+X_{3},$ (S57)
$X_{2}$ $\to X_{1}+X_{3},$ (S58)
$X_{3}$ $\to X_{1}+X_{2}.$ (S59)
Observe that the last network is exactly the type V autocatalytic core.
S0.7 E4. $\tilde{\Lambda}$ vs $\delta$ for $M^{+}=M^{-}=10$
The figure below shows that the spectrum of values of $\alpha$ becomes continuous in the limit that $M^{+}$ and $M^{-}$ become large.
Figure S2: Growth parameter $\delta=\alpha-1$ versus normalized growth rate $\tilde{\Lambda}=\Lambda/\|\mathbb{S}^{-}\|_{k}$ for random autonomous, unambiguous CRNs with $2\leq|\mathcal{S}|,|\mathcal{R}|\leq 12$ with $M^{+}=M^{-}=10$. Left: autocatalytic networks ($\alpha>1$, $\delta>0$) and networks which admit a steady-state flux ($\alpha=1$, $\delta=0$). Right: autoinhibitory networks ($\alpha<1$, $\delta<0$).